OpenStudy (anonymous):

The length of a rectangle is 3 feet longer than its width. If its length is decreased by 3 feet and its width is increased by 2 feet, its area would be decreased by 8 square feet. What is the length of the rectangle?

OpenStudy (y2o2):

l -w = 3 (l-3)(w+2) + 8 = lw so l = 11 feet and w = 8 feet

OpenStudy (anonymous):

Thank you... I was trying to do a totally different equation thats why it didnt work ha ha thanks for helping me!

OpenStudy (y2o2):

Don't worry , we all o so ;)